# Moduli spaces of untwisted wild Riemann surfaces

### v1

@@ 2.4.3. Definition of $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p,\underline{\mathbf{d}}}$
Let us verify that the properties defining a stack (cf. section 4.1.4) are satisfied.
@@
    Objects glue. Let $(B_{i})_{i\in I}$ be an open cover of a manifold $B$. Suppose we are given objects $(\pi_{i}\colon\Sigma_{i}\rightarrow B_{i},\underline{a}_{i},\underline{Q}_{i})$ of $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}(B_{i})$ for each $i$, together with isomorphisms of their restrictions to $B_{i}\cap B_{j}$ satisfying the cocycle condition on triple intersections. We need to show that there is $(\pi\colon\Sigma\rightarrow B,\underline{a},\underline{Q})$ pulling back to $(\pi_{i}\colon\Sigma_{i}\rightarrow B_{i},\underline{a}_{i},\underline{Q}_{i})$ over each $B_{i}$. Since $\mathsf{M}_{g,m}$ is a stack, we do get $(\pi\colon\Sigma\rightarrow B,\underline{a})$ with the desired property. Now, each irregular irregular type $\underline{Q}_{i}$ is a collection of sections of the pullbacks to $B_{i}$ of $\mathfrak{t}\otimes(\pi_{*}\mathscr{T}_{\Sigma,a_{i}}^{\leq p}),1\leq i\leq m$. Such sections glue to a section on $B$ because $\pi_{*}\mathscr{T}_{\Sigma,a_{i}}^{\leq p}$ is a sheaf on $B$.
@@
    Isomorphisms glue. Isomorphisms of objects in $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}(B)$ are by definition isomorphisms of Riemann surfaces over $B$ (respecting the extra structures), which glue. ∎
We can now state the main result of this section.
2.4.6 Theorem.
Given integers $g\geq 0,m\geq 1,p\geq 1$, and given $\underline{\mathbf{d}}=(d_{\alpha,i})_{1\leq i\leq m,\alpha\in\Phi}$ with $0\leq d_{\alpha,i}\leq p$, the following assertions hold true.
@@
    The stack $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}$ is an analytic stack.
@@
    The map $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}\rightarrow\mathsf{M}_{g,m}$ is representable by vector bundles. More precisely, let $B$ be a complex manifold and $B\rightarrow\mathsf{M}_{g,m}$ a map corresponding to an $m$-pointed genus $g$ Riemann surface $(\Sigma\rightarrow B,\underline{a})$, where $\underline{a}=(a_{1},\ldots,a_{m})$. The fibre product $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq p}\times_{\mathsf{M}_{g,m}}B$ is isomorphic to $\mathscr{I\!\!T}^{\mathfrak{t},\leq p}_{B,a_{1},\ldots,a_{m}}$.

### v2

@@ 2.4.3. Definition of $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}$
Let us verify that the properties defining a stack (cf. section A.1.4) are satisfied.
Objects glue:
Let $(B_{j})_{j\in J}$ be an open cover of a manifold $B$. Suppose we are given objects $(\pi_{j}\colon\Sigma_{j}\rightarrow B_{j},\underline{a}_{j},\underline{Q}_{j})$ of $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}(B_{j})$ for each $j$, together with isomorphisms of their restrictions to $B_{j}\cap B_{j^{\prime}}$ satisfying the cocycle condition on triple intersections. We need to show that there is $(\pi\colon\Sigma\rightarrow B,\underline{a},\underline{Q})$ pulling back to $(\pi_{j}\colon\Sigma_{j}\rightarrow B_{j},\underline{a}_{j},\underline{Q}_{j})$ over each $B_{j}$. Since $\mathsf{M}_{g,m}$ is a stack, we do get $(\pi\colon\Sigma\rightarrow B,\underline{a})$ with the desired property. Now, each $\underline{Q}_{j}$ is a collection of sections over $B_{j}$ of $\mathfrak{t}\otimes_{\mathbf{C}}(\pi_{*}\mathscr{T}_{\Sigma,a_{i}}^{\leq p_{i}}),$ for $1\leq i\leq m$. Such sections glue to a section on $B$ because $\pi_{*}\mathscr{T}_{\Sigma,a_{i}}^{\leq p_{i}}$ is a sheaf on $B$.
Isomorphisms glue:
Isomorphisms of objects in $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}(B)$ are by definition isomorphisms of Riemann surfaces over $B$ (respecting the extra structures), which glue. ∎
We can now state the main result of this section. See Definition A.1.8 for the definition of analytic stack.
2.4.6 Theorem.
Given integers $g\geq 0,m\geq 1$, and tuples
$\underline{p}=(p_{1},\ldots,p_{m})\in\mathbf{Z}_{\geq 0}^{m},\quad\underline{\mathbf{d}}=(d_{\alpha,i})_{\alpha\in\Phi,1\leq i\leq m},\qquad\text{ with }0\leq d_{\alpha,i}\leq p_{i},$
the following assertions hold true.
@@
    The stack $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}$ is analytic.
@@
    The map $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}\rightarrow\mathsf{M}_{g,m}$ is representable by vector bundles. More precisely, let $B$ be a complex manifold and $B\rightarrow\mathsf{M}_{g,m}$ a map corresponding to an $m$-pointed genus $g$ Riemann surface $(\Sigma\rightarrow B,\underline{a})$, where $\underline{a}=(a_{1},\ldots,a_{m})$. The fibre product $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p}}\times_{\mathsf{M}_{g,m}}B$ is isomorphic to $\mathscr{I\!\!T}^{\mathfrak{t},\leq\underline{p}}_{B,\underline{a}}$.
@@
    The assignment $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}\colon\mathrm{Man}_{\mathbf{C}}\rightarrow\mathrm{Groupoids}$ sending $B$ to $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}(B)$ is an analytic stack.
@@
    The map $\mathsf{WM}_{g,m}^{\mathfrak{t},\leq\underline{p},\underline{\mathbf{d}}}\rightarrow\mathsf{M}_{g,m}$ is representable by manifolds which are locally products of hyperplane complements in affine spaces (in particular, it is a submersion).